# Beam Cross Sections Create Mixtures: Improving Feature Localization in Secondary Electron Imaging

### v1

@@ V-B Edge Location Estimates
Interpolation: The current practice of SE imaging is based on computing the baseline estimate (28) at each pixel prior to any subsequent processing. In the absence of an explicit model for the effect of the beam cross section, one could declare an edge location to be where the estimated SE yield crosses a threshold value. For example, the edge location could be estimated at the location where the linearly interpolated estimated SE yield equals $(\eta_{1}+\eta_{2})/2$. In our specific case of left-to-right scanning from lower SE yield to higher SE yield, as illustrated in Fig. 2(b), the interpolated SE yield estimate function may cross the threshold multiple times. When this happens, we use the sample mean of all the upward crossings.
@@
$\mathcal{L}_{\mathrm{scan}}(\gamma)=\prod_{k=0}^{\ell-1}\mathop{\mathrm{Poisson}}(\widetilde{m}_{k}\,;\,\lambda\rho_{k})\prod_{i=1}^{\widetilde{m}_{k}}\mathrm{P}_{\widetilde{X}}(\widetilde{x}_{i,k}\,;\,q_{k},\eta_{1},\eta_{2}),$ (70)
@@
$\widehat{\gamma}_{\mathrm{MLE}}=\operatorname*{arg\,max}_{\gamma}\mathcal{L}_{\mathrm{scan}}(\gamma).$ (71)
Mismatched MLE: To isolate the effect of the mixture model from the effect of TRM, we also consider an estimator that maximizes the likelihood for Poisson-distributed data; i.e., the estimator uses the conventional convolutional model for the beam cross-section rather than the mixture model. As above, for any candidate edge location $\gamma$ and each horizontal raster scan location $k$, we have a corresponding parameter $q_{k}$. This implies a mean $\mu_{k}$ given by (10), and each SE count distribution $X_{i,k}$ is assumed to be $\mathop{\mathrm{Poisson}}(\mu_{k})$ as in (12). Differing from above, the thinning parameter in the distribution of $\widetilde{M}$ is $\rho_{k}^{\prime}=1-e^{-\mu_{k}}$. Also differing from above, each $\widetilde{X}_{i,k}$ has a zero-truncated Poisson distribution with parameter $\mu_{k}$, $\mathop{\mathrm{ZTPoisson}}(\mu_{k})$. Combining the likelihoods over the scan positions gives
$\mathcal{L}_{\mathrm{conv}}(\gamma)=\prod_{k=0}^{\ell-1}\mathop{\mathrm{Poisson}}(\widetilde{m}_{k}\,;\,\lambda\rho_{k}^{\prime})\prod_{i=1}^{\widetilde{m}_{k}}\mathop{\mathrm{ZTPoisson}}(\widetilde{x}_{i,k}\,;\,\mu_{k}),$ (72)
where $\rho_{k}^{\prime}$ and $\mu_{k}$ are functions of $\gamma$, $k$, $\eta_{1}$, $\eta_{2}$, and $\sigma_{\mathrm{b}}$. Maximizing this function gives the mismatched MLE:
$\widehat{\gamma}_{\mathrm{MMLE}}=\operatorname*{arg\,max}_{\gamma}\mathcal{L}_{\mathrm{conv}}(\gamma).$ (73)
@@ V-C Simulations of Edge Localization
We performed Monte Carlo simulations to evaluate the performances of the three edge estimators, with the MLE and MMLE implemented by grid search with a grid spacing of 0.01. Each Monte Carlo trial has pseudorandom data generated according to the model in Section II on a one-dimensional scan grid of $\{0,1,\ldots,99\}$ (i.e., $\ell=100$) Detector noise is not modeled, so a simulated TRM dataset is $\{(\widetilde{m}_{k},\{\widetilde{x}_{i,k}\}_{i=1}^{\widetilde{m}_{k}})\}_{k=0}^{\ell-1}$. Simulation parameters include the edge location $\gamma$, beam width $\sigma_{\mathrm{b}}$, dose $\lambda$, and SE yields on each side of the edge, $\eta_{1}$ and $\eta_{2}$. Where applicable, the estimators use $\sigma_{\mathrm{b}}$, $\lambda$, $\eta_{1}$, and $\eta_{2}$; only $\gamma$ is unknown. Sets of simulations vary one parameter while keeping others fixed. Each reported result is based on 300 Monte Carlo trials, which allows for computation of (estimated) bias, (sample) standard deviation, and (estimated) root mean-squared error (RMSE). The error bars in the bias plots represent $\pm 1$ standard deviation of the sample mean. We similarly computed $\pm 1$ standard deviation interval endpoints for the sample variance and the sample MSE. The marked error bars in the standard deviation and RMSE plots use the square roots of these quantities. The reciprocal of the Fisher information (68) is the Cramér–Rao lower bound on the MSE of any unbiased estimate of $\gamma$ from a TRM dataset. The square root of this CRB is included in all RMSE plots.
@@
The MLE has very low bias, and its standard deviation and RMSE follow the $\lambda^{-1/2}$ dependence on $\lambda$ predicted by $\mathcal{I}_{\mathrm{scan}}(\gamma\,;\,\eta_{1},\eta_{2},\lambda,\sigma_{\mathrm{b}})$ in (68) being linear in $\lambda$. Along with the $\lambda^{-1/2}$ scaling behavior, the quantitative match of the RMSE to the square root of the CRB is good here and in most subsequent simulation results, showing that the MLE is approximately efficient.
@@
The interpolation estimator has very low bias at high doses—but significant bias at low doses—and higher standard deviation than the MLE and MMLE. The bias being positive and vanishing at high enough doses can be explained as follows. Gross inaccuracy of the interpolation estimator is most likely when there are multiple crossings of the threshold $(\eta_{1}+\eta_{2})/2=5.5$. Far to the left of the edge, $\widehat{\eta}_{\rm baseline}$ has mean $\eta_{1}=1$ and variance $(\eta_{1}+\eta_{1}^{2})/\lambda=2/\lambda$ (see (38) and (39)). Thus, even at the lowest shown dose of $\lambda=20$, spurious threshold crossings (i.e., $\widehat{\eta}_{\rm baseline}>5.5$) are very unlikely. Conversely, far to the right of the edge, $\widehat{\eta}_{\rm baseline}$ has mean $\eta_{2}=10$ and variance $(\eta_{2}+\eta_{2}^{2})/\lambda=110/\lambda$, and spurious threshold crossings (i.e., $\widehat{\eta}_{\rm baseline}<5.5$) happen with nonnegligible probability unless $\lambda$ is large enough.
For panel (b), we choose $\lambda=200$ so that the dose is high enough to avoid poor performance of interpolation caused by high likelihood of multiple threshold crossings. We vary $\sigma_{\mathrm{b}}$ from 0.1 to 3 while keeping $\gamma=50.2$, $\eta_{1}=1$, and $\eta_{2}=10$. The RMSE comparison again shows the superiority of the MLE over the MMLE and interpolation. The signs of biases match the results in panel (a), with MLE having very low bias, MMLE having negative bias, and interpolation having positive bias. For MMLE, as $\sigma_{\mathrm{b}}$ shrinks the bias shrinks because the model is substantially mismatched only at raster scan locations where the mixture parameter is close to neither 0 nor 1. For interpolation, since the dose is relative large, the source of bias is different than in panel (a). As $\sigma_{\mathrm{b}}$ shrinks, the beam interaction with the sample edge vanishes unless the edge is located very close to a raster scan position. For $\gamma=50.2$, when $\sigma_{\mathrm{b}}$ is very small and $\lambda$ is sufficiently high, the interpolation estimate should usually be 50.5 because $\widehat{\eta}_{\rm baseline}\approx 1$ at $k=50$ and $\widehat{\eta}_{\rm baseline}\approx 10$ at $k=51$, with no spurious threshold crossings. Thus, the bias is 0.3. This phenomenon is also exhibited by the monotonic increase of the standard deviation of the interpolation estimator as $\sigma_{\mathrm{b}}$ is increased. We comment on the dependence of the bias on $\gamma$ in discussing panel (c).
@@
Panel (c) differs from panel (b) only in changing the edge location to $\gamma=50.5$. One striking change is to the bias of interpolation at low values of $\sigma_{\mathrm{b}}$. This bias is greatly reduced because $\widehat{\eta}_{\rm baseline}\approx 1$ at scan position 50 combined with $\widehat{\eta}_{\rm baseline}\approx 10$ at scan position 51 gives an interpolation estimate near 50.5 by happenstance; for interpolation with small $\sigma_{\mathrm{b}}$, the edge location being at a midpoint between neighboring scan locations is highly favorable. The bias of the MLE is not significantly changed, but its standard deviation is significantly increased at low values of $\sigma_{\mathrm{b}}$. Referring to Fig. 7(a), $\gamma=50.5$ is a near minimizer of the FI about $\gamma$ in the scan data. Indeed, the minima of the RMSE and CRB shown in panel (c) occur near the $\sigma_{\mathrm{b}}^{*}=0.33$ value found in Fig. 7(b).
Finally, in panel (d) we vary the right-side SE yield $\eta_{2}$ from 2 to 20 while fixing $\eta_{1}=1$, $\gamma=50.2$, $\sigma_{\mathrm{b}}=1$, and $\lambda=200$. As before, the MLE has a very low bias, and its standard deviation and RMSE are the lowest across all values of $\eta_{2}$. The root causes of biases of the interpolation estimator and MMLE are as discussed in regard to panel (a). For interpolation, the bias is most pronounced when the gap between $\eta_{1}$ and $\eta_{2}$ is small, making spurious threshold crossings relatively more likely. For MMLE, the bias is smallest where the excess variance of the mixture is smallest.
A consistent trend across all experiments is that the MLE estimate is unbiased and has the lowest standard deviation among all estimators. It thus has the best performance for varied sample and imaging parameters. The RMSE in many cases is well below the unit grid spacing, showing the possibility of super-resolution localization.
@@ V-D Experimental Edge Localization
We performed proof-of-concept experiments with real data to test estimator performances. We used a Zeiss Orion Nanolab HIM to image a sample of gold nanofabricated on a silicon substrate. This sample and the imaging conditions are described in prior work [27], from which we use SE yield values of $\eta_{1}=1.82$ for silicon and $\eta_{2}=2.75$ for gold. We created three datasets, each consisting of 51 scan lines with 100 raster scan locations across an edge, analogous to the Monte Carlo simulations, with a dose per scan location of $\lambda=20$. The dataset generation process is described in more detail in the supplement.

### v2

@@ V-B Edge Location Estimates
Interpolation: The current practice of SE imaging is based on computing the baseline estimate (25) at each pixel prior to any subsequent processing. If the knowledge of $\eta_{1}$ and $\eta_{2}$ is available a priori, one can clip this $\widehat{\eta}_{\rm baseline}$ to lie in $\eta_{1},\eta_{2}]$. In the absence of an explicit model for the effect of the beam cross section, one could declare an edge location to be where the estimated SE yield crosses a threshold value. For example, the edge location could be estimated at the location where the linearly interpolated estimated SE yield equals $(\eta_{1}+\eta_{2})/2$. We refer to the edge estimator based on level crossings of $\widehat{\eta}_{\rm baseline}$ as the Interpolation estimator, while the estimator that uses the clipped version of $\widehat{\eta}_{\rm baseline}$ as the Interpolation-clip estimator. In our specific case of left-to-right scanning from lower SE yield to higher SE yield, as illustrated in Fig. 2(b), the interpolated SE yield estimate function may cross the threshold multiple times. When this happens, we use the sample mean of all the upward crossings.
@@
$\mathcal{L}_{\mathrm{scan}}^{\mathrm{mix}}(\gamma)=\prod_{k=0}^{\ell-1}\mathop{\mathrm{Poisson}}(\widetilde{m}_{k}\,;\,\lambda\rho_{k})\prod_{i=1}^{\widetilde{m}_{k}}\mathrm{P}_{\widetilde{X}}(\widetilde{x}_{i,k}\,;\,q_{k},\eta_{1},\eta_{2}),$ (64)
@@
$\widehat{\gamma}_{\mathrm{MLE}}=\operatorname*{arg\,max}_{\gamma}\mathcal{L}_{\mathrm{scan}}^{\mathrm{mix}}(\gamma).$ (65)
Mismatched MLE: To isolate the effect of the mixture model from the effect of TRM, we also consider an estimator that maximizes the likelihood for Poisson-distributed data; i.e., the estimator treats the effect of the beam cross-section as a convolution. As above, for any candidate edge location $\gamma$ and each horizontal raster scan location $k$, we have a corresponding parameter $q_{k}$ from (6). This implies a mean $\mu_{k}$ given by (9), and each SE count distribution $X_{i,k}$ is assumed to be $\mathop{\mathrm{Poisson}}(\mu_{k})$. Differing from above, the thinning parameter in the distribution of $\widetilde{M}$ is $\rho_{k}^{\prime}=1-e^{-\mu_{k}}$. Also differing from above, each $\widetilde{X}_{i,k}$ has a zero-truncated Poisson distribution with parameter $\mu_{k}$, $\mathop{\mathrm{ZTPoisson}}(\mu_{k})$. Combining the likelihoods over the scan positions gives
$\mathcal{L}_{\mathrm{scan}}^{\mathrm{Poisson}}$ $(\gamma)$
$=\prod_{k=0}^{\ell-1}\mathop{\mathrm{Poisson}}(\widetilde{m}_{k}\,;\,\lambda\rho_{k}^{\prime})\prod_{i=1}^{\widetilde{m}_{k}}\mathop{\mathrm{ZTPoisson}}(\widetilde{x}_{i,k}\,;\,\mu_{k}),$
where $\rho_{k}^{\prime}$ and $\mu_{k}$ are functions of $\gamma$, $k$, $\eta_{1}$, $\eta_{2}$, and $\sigma_{\mathrm{b}}$. Maximizing this function gives the mismatched MLE (MMLE):
$\widehat{\gamma}_{\mathrm{MMLE}}=\operatorname*{arg\,max}_{\gamma}\mathcal{L}_{\mathrm{scan}}^{\mathrm{Poisson}}(\gamma).$ (66)
@@ V-C Simulations of Edge Localization
We performed Monte Carlo simulations to evaluate the performances of the four edge estimators, with the MLE and MMLE implemented by grid search with a grid spacing of 0.01. Each Monte Carlo trial has pseudorandom data generated according to the model in Section II on a one-dimensional scan grid of $\{0,1,\ldots,99\}$ (i.e., $\ell=100$). Detector noise is not modeled, so a simulated TRM dataset is $\{(\widetilde{m}_{k},\{\widetilde{x}_{i,k}\}_{i=1}^{\widetilde{m}_{k}})\}_{k=0}^{\ell-1}$. Simulation parameters include the edge location $\gamma$, beam width $\sigma_{\mathrm{b}}$, dose $\lambda$, and SE yields on each side of the edge, $\eta_{1}$ and $\eta_{2}$. Where applicable, the estimators use $\sigma_{\mathrm{b}}$, $\lambda$, $\eta_{1}$, and $\eta_{2}$; only $\gamma$ is unknown. In the stylized application considered here, $\eta_{1}$ and $\eta_{2}$ could be estimated well by aggregating data from many pixels far from the edge. Sets of simulations vary one parameter while keeping others fixed. Each reported result is based on 300 Monte Carlo trials, which allows for computation of (estimated) bias, (sample) standard deviation, and (estimated) root mean-squared error (RMSE). The error bars in the bias plots represent $\pm 1$ standard deviation of the sample mean. We similarly computed $\pm 1$ standard deviation interval endpoints for the sample variance and the sample MSE. The marked error bars in the standard deviation and RMSE plots use the square roots of these quantities. The reciprocal of the Fisher information (62) is the Cramér–Rao lower bound on the MSE of any unbiased estimate of $\gamma$ from a TRM dataset. The square root of this CRB is included in all RMSE plots.
@@
The MLE has very low bias, and its standard deviation and RMSE follow the $\lambda^{-1/2}$ dependence on $\lambda$ predicted by $\mathcal{I}_{\mathrm{scan}}(\gamma\,;\,\eta_{1},\eta_{2},\lambda,\sigma_{\mathrm{b}})$ in (62) being linear in $\lambda$. Along with the $\lambda^{-1/2}$ scaling behavior, the quantitative match of the RMSE to the square root of the CRB is good here and in most subsequent simulation results.
@@
Here we observe that the interpolation estimators coincide and clipping does not give any advantage. Both estimators have very low bias at high doses—but significant bias at low doses—and higher standard deviation than the MLE and MMLE. The bias being positive and vanishing at high enough doses can be explained as follows. Gross inaccuracy of the interpolation estimator is most likely when there are multiple crossings of the threshold $(\eta_{1}+\eta_{2})/2=5.5$. Far to the left of the edge, $\widehat{\eta}_{\rm baseline}$ has mean $\eta_{1}=1$ and variance $(\eta_{1}+\eta_{1}^{2})/\lambda=2/\lambda$ (see (33) and (34)). Thus, even at the lowest shown dose of $\lambda=20$, spurious threshold crossings (i.e., $\widehat{\eta}_{\rm baseline}>5.5$) are very unlikely. Conversely, far to the right of the edge, $\widehat{\eta}_{\rm baseline}$ has mean $\eta_{2}=10$ and variance $(\eta_{2}+\eta_{2}^{2})/\lambda=110/\lambda$, and spurious threshold crossings (i.e., $\widehat{\eta}_{\rm baseline}<5.5$) happen with nonnegligible probability unless $\lambda$ is large enough.
For panel (b), we choose $\lambda=200$ so that the dose is high enough to avoid poor performance of interpolation caused by high likelihood of multiple threshold crossings. We vary $\sigma_{\mathrm{b}}$ from 0.1 to 3 while keeping $\gamma=50.2$, $\eta_{1}=1$, and $\eta_{2}=10$. The RMSE comparison again shows the superiority of the MLE over the MMLE and interpolation. The signs of biases match the results in panel (a), with MLE having very low bias, MMLE having negative bias, and interpolation having positive bias. For MMLE, as $\sigma_{\mathrm{b}}$ shrinks the bias shrinks because the model is substantially mismatched only at raster scan locations where the mixture parameter is close to neither 0 nor 1. Here we first see a performance difference between the interpolation estimators, with clipping leading to noticeably lower standard deviation values for the lower values of $\sigma_{\mathrm{b}}$. The biases of the interpolation estimators are nearly equal, and bias dominates the RMSE at low $\sigma_{\mathrm{b}}$ values, so there is no noticeable difference in the RMSEs of the interpolation estimators. For both interpolation estimators, since the dose is relatively large, the source of bias is different than in panel (a). As $\sigma_{\mathrm{b}}$ shrinks, the beam interaction with the sample edge vanishes unless the edge is located very close to a raster scan position. For $\gamma=50.2$, when $\sigma_{\mathrm{b}}$ is very small and $\lambda$ is sufficiently high, the interpolation estimate should usually be 50.5 because $\widehat{\eta}_{\rm baseline}\approx 1$ at $k=50$ and $\widehat{\eta}_{\rm baseline}\approx 10$ at $k=51$, with no spurious threshold crossings. Thus, the bias is 0.3. This phenomenon is also exhibited by the monotonic increase of the standard deviation of the interpolation estimator as $\sigma_{\mathrm{b}}$ is increased. We comment on the dependence of the bias on $\gamma$ in discussing panel (c).
@@
Panel (c) differs from panel (b) only in changing the edge location to $\gamma=50.5$. One striking change is to the bias of the interpolation methods at low values of $\sigma_{\mathrm{b}}$. This bias is greatly reduced because $\widehat{\eta}_{\rm baseline}\approx 1$ at scan position 50 combined with $\widehat{\eta}_{\rm baseline}\approx 10$ at scan position 51 gives an interpolation estimate near 50.5 by happenstance; for interpolation with small $\sigma_{\mathrm{b}}$, the edge location being at a midpoint between neighboring scan locations is highly favorable. In this case, clipping of $\widehat{\eta}_{\rm baseline}$ gives a slight improvement in standard deviation and RMSE at lower values of $\sigma_{\mathrm{b}}$. The bias of the MLE is not significantly changed, but its standard deviation is significantly increased at low values of $\sigma_{\mathrm{b}}$. Referring to Fig. 7(a), $\gamma=50.5$ is a near minimizer of the FI about $\gamma$ in the scan data. Indeed, the minima of the RMSE and CRB shown in panel (c) occur near the $\sigma_{\mathrm{b}}^{\star}=0.33$ value found in Fig. 7(b).
The deviation of MLE performance from the CRB at $\sigma_{\mathrm{b}}=0.1$ warrants further study. There is no contradiction in the low exhibited RMSE because the MLE is not unbiased at this very small $\sigma_{\mathrm{b}}$. At the two scan locations $k$ closest to the edge (50 and 51), the normalized distance $|k-\gamma|/\sigma_{\mathrm{b}}$ between the scan location and the edge at $\gamma=50.5$ is 5, implying $\min\{q,1-q\}\approx 3\times 10^{-7}$. Incident particles are very unlikely to strike on opposite side of the edge relative to the nominal scan location, so the mixtures are nearly indistinguishable from (single-component) Poisson distributions.
Finally, in panel (d) we vary the right-side SE yield $\eta_{2}$ from 2 to 20 while fixing $\eta_{1}=1$, $\gamma=50.2$, $\sigma_{\mathrm{b}}=1$, and $\lambda=200$. As before, the MLE has a very low bias, and its standard deviation and RMSE are the lowest across all values of $\eta_{2}$. The root causes of biases of the interpolation estimator and MMLE are as discussed in regard to panel (a). Here also we observe that the two interpolation estimators coincide; clipping does not give any advantage. For both the interpolation estimators, the bias is most pronounced when the gap between $\eta_{1}$ and $\eta_{2}$ is small, making spurious threshold crossings relatively more likely. For MMLE, the bias is smallest where the excess variance of the mixture is smallest. A trend across most experiments is that the MLE is approximately unbiased and has the lowest standard deviation among all estimators. It thus has the best performance for varied sample and imaging parameters. The RMSE in many cases is well below the unit grid spacing, showing the possibility of sub-pixel localization.
@@ V-D Experimental Edge Localization
We performed proof-of-concept experiments with real data to test estimator performances. We used a Zeiss Orion Nanolab HIM to image a sample of gold nanofabricated on a silicon substrate. This sample and the imaging conditions are described in prior work [27], from which we use SE yield values of $\eta_{1}=1.82$ for silicon and $\eta_{2}=2.75$ for gold. We created three datasets, each consisting of 51 scan lines with 100 raster scan locations across an edge, analogous to the Monte Carlo simulations, with a dose per scan location of $\lambda=20$. The dataset generation process is described in more detail in the Supplementary Note 4.
Real HIM datasets have several complexities, such as the conversion of SE counts to detector voltages [45] and increase in SE yield at edges [29, 46, 47, 6, 49, 48]. Furthermore, lack of availability of $\sigma_{\mathrm{b}}$ and the ground truth edge location make direct application and evaluation of the methods described here challenging. We converted detector voltages to SE counts using a hard thresholding based on prior work [27], and we used a defocused beam to reduce edge brightening. Since $\sigma_{\mathrm{b}}$ is not otherwise available, we computed a joint MLE from all the available data (all lines of the three datasets) of the beam width and collection of edge locations. To generate proxies for the ground truth edge location $\gamma$, we applied the interpolation estimator to an emulation of the very high dose of $\lambda=1020$ by concatenating the data from 51 line scans. As we observe in Fig. 8(a), the interpolation estimate has very low bias and variance as the dose increases, making it a good ground truth proxy.
The results of running the estimators on experimental data are shown in Table I. We observed that both the interpolation methods gave the same performance because clipping never influenced a $\gamma$ estimate. Hence these estimators are not shown separately in the table. The MLE has lowest bias across all the three datasets, followed by the MMLE and interpolation methods. Both MLE and MMLE have lower standard deviations than the interpolation methods, similar to the simulation results. MLE consistently achieves the lowest RMSE across all three datasets, outperforming the interpolation method by a factor of 4 to 7 and surpassing MMLE by about 1.5 times. These results align with the trends observed in our simulation experiments.
TABLE I: Experimental results with three HIM datasets